OpenStudy (anonymous):

Someone wanna help me out? 2^(4x+2)=8^(x-3)

OpenStudy (barrycarter):

Hint: logarithms

OpenStudy (zzr0ck3r):

log two of both sides gives 4x+2 = (x-3)log_2(8) so 4x+2 = 3(x-3)

OpenStudy (zzr0ck3r):

do you see why?

OpenStudy (zzr0ck3r):

when I write log I mean log base 2 log(2^(4x+2)) = (4x+2)log(2) = 4x+2

OpenStudy (anonymous):

-_- Im retarded. Yeah, I got -11

OpenStudy (zzr0ck3r):

and on the rhs log(8^(x-3)) = (x-3)log(8) now 2^? = 8 3 so 3(x-3)

OpenStudy (zzr0ck3r):

correct

OpenStudy (anonymous):

Thanks!

OpenStudy (zzr0ck3r):

np